Calculate the pH of a 0.021 M NaCN solution. [Kb(NaCN) = 5.0 ´ 10-5]
1
Expert's answer
2020-10-19T14:26:35-0400

Since , NaCN is formed by a mixture of a strong base and weak acid we need Ka of HCN to calculate the pH

I am just solving it using the Kb

Ph = 7+0.5Pkb-0.5logC

Ph= 7 - 0.5logKb -0.5log0.021

Ph =7 + 2.15 +0.83

= 9.98